# Impact of QCD Energy Evolution on Observables in Heavy-Ion Collisions

### v1

@@ III Space-time evolution
As mentioned above, the gluon fields after the collision are obtained in the standard way Krasnitz:1998ns; Krasnitz:1999wc; Krasnitz:2000gz; Schenke:2012wb from the Wilson lines determined with one of the above prescriptions. Further, the early-time evolution in the Glasma phase up to $\tau_{0}=0.4$ fm/c is obtained by solving the source-free Yang-Mills equations as implemented in the IP-Glasma model. The energy-momentum tensor $T^{\mu\nu}$ of the system is computed from the gluon fields at $\tau=\tau_{0}$, providing the initial condition for relativistic viscous hydrodynamic simulations, which are performed using MUSIC Schenke:2010nt; Schenke:2010rr; Paquet:2015lta.
@@
Since the strong coupling constant factors out of the classical Yang-Mills evolution equations (see discussion in Schenke:2020mbo), we determine an overall normalization factor $K$ for the energy-momentum tensor $T^{\mu\nu}$ by matching to the measured charged hadron multiplicity in the 0–5% most central Au+Au collisions at $\sqrt{s_{\mathrm{NN}}}=200$ GeV. Once fixed, the JIMWLK evolution predicts the energy dependence of particle production across different collision energies. This stands in contrast to the conventional Glauber-based approach, where the normalization is typically re-tuned separately for each center-of-mass energy. We note that running coupling effects can be expected to result in $K$ having a logarithmic dependence on the center-of-mass energy. As running coupling corrections are included in the JIMWLK evolution employed in this work, analogous corrections could, in principle, be applied to the $K$ factor. This would introduce an additional dependence on the choice of running coupling scheme and is therefore left for future work.
@@
Focusing exclusively on midrapidity observables, we assume boost invariance in the hydrodynamic evolution. We employ an equation of state derived from lattice QCD calculations Bazavov:2014pvz; Moreland:2015dvc, and include both shear and bulk viscous effects. These are treated using the Denicol-Niemi-Molnar-Rischke (DNMR) second-order viscous hydrodynamic formalism Denicol:2012cn, which incorporates spatial gradient corrections up to second order.
@@
The transition from the fluid to hadronic degrees of freedom is performed on an isothermal hypersurface defined by a constant energy density $e_{\mathrm{sw}}=0.18~\mathrm{GeV/fm}^{3}$, using the Cooper-Frye particlization procedure with viscous corrections implemented via Grad’s 14-moment method Huovinen:2012is; Shen:2014vra; Zhao:2022ugy. The resulting hadrons are subsequently evolved in the dilute hadronic phase using the UrQMD transport model Bass:1998ca; Bleicher:1999xi, which accounts for hadronic scatterings and decays.
@@
Here $s$ and $T$ are the entropy density and temperature, respectively. We tune the parameters to reproduce experimental data on charged hadron multiplicity PHENIX:2003iij, mean transverse momentum STAR:2008med, and azimuthal flow coefficients STAR:2015mki measured in Au+Au collisions at $\sqrt{s_{\mathrm{NN}}}=200$ GeV. The IP-Glasma+JIMWLK setup is used for this tune. The purpose of this work is to quantify the effects of including the small-$x$ JIMWLK evolution in the initial state description. Consequently, the transport properties are constrained using the RHIC data only, as there is only moderate evolution from initial $x_{0}=0.01$ to RHIC kinematics corresponding to $x=0.0025$. The obtained parameters are $\eta_{0}=0.14$, $T_{0}=0.18$ GeV, b=4b=-4 GeV-1, $(\zeta/s)_{\max}=0.123$, $\lambda_{\zeta}=-0.12$, and $w_{\zeta}=0.05$ GeV.
@@ IV.1 Heavy ion collisions
Figure 1: Charged hadron multiplicity distributions in Pb+Pb collisions at the LHC 5.02 TeV and 5.36 TeV, and Au+Au collisions at RHIC 200 GeV energies. The experimental data are taken from ALICE:2015juo; ALICE:2025cjn; PHENIX:2004vdg.
@@
Both prescriptions for the initial state energy dependence reproduce the multiplicity in the most central collisions. On the other hand, in more peripheral Pb+Pb collisions beyond a centrality of $\sim 20\%$, inclusion of the JIMWLK evolution results in significantly flatter multiplicity distributions. We interpret this to be due to the geometry evolution included in the “+JIMWLK” setup, which modifies the overlap by blurring the edges of the incoming nuclei, and consequently changes the shape of the multiplicity distribution with centrality. This geometry evolution has been previously studied e.g. in Refs. Schlichting:2014ipa; Mantysaari:2018zdd; Mantysaari:2022sux. We have checked that the differences in the viscous entropy production during the late-time hydrodynamic phase are not significant.
@@
Figure 2: Anisotropic flow $v_{n}$ in Au+Au collisions 200 GeV. The experimental data are from STAR:2016vqt; STAR:2017idk; STAR:2019zaf.
@@
Figure 3: Collective flow $v_{n}$ in Pb + Pb collisions 5.02 TeV. The experimental data are from ALICE:2016ccg.
@@
Figure 4: The $v_{2}\{4\}/v_{2}\{2\}$ ratio in Pb + Pb collisions at 5.02 TeV. The experimental data are from ALICE:2016ccg.
@@
The effect of JIMWLK evolution on anisotropic flow coefficients was not studied in the companion letter Mantysaari:2025tcg. To set the stage, we first show our results for $v_{2}\{2\},v_{3}\{2\}$ and $v_{4}\{2\}$ in Au+Au collisions at 200 GeV center-of-mass energy in Fig. 2. Note that shear and bulk viscosities have been adjusted to describe this data in the “+JIMWLK” setup, and kept the same in the $Q_{s}(x)$ setup - see discussion in Sec. III. There is only moderate evolution from the initial $x=0.01$ to the typical midrapidity value at RHIC, $x=0.0025$. Yet, there is some difference between our results for $v_{2}\{2\}$ and $v_{2}\{4\}$ for centralities larger than 40% as well as for $v_{3}\{2\}$ for all centralities. In all cases the +JIMWLK result is lower than the corresponding $Q_{s}(x)$ result. This can again be explained by the smoothing effect of the JIMWLK evolution on the initial state geometry, which reduces the initial eccentricities, as we have checked explicitly.
@@
Similar trends are seen for Pb+Pb collisions at 5.02 TeV: The centrality dependence of the flow harmonics $v_{2}\{2\},v_{3}\{2\}$ and $v_{4}\{2\}$ is shown in Fig. 3. Again, the JIMWLK evolution is found to suppress all flow harmonics compared to the $Q_{s}(x)$ setup. Interestingly, the differences are similar to those observed in the lower energy Au+Au collisions, suggesting that the dominant effect from the small-$x$ evolution is the smoothening of the initially lumpy nuclear substructure early in the evolution. Faster evolution at larger $x$ is expected as saturation effects that slow down the evolution become more prominent as $x$ decreases. The ALICE data ALICE:2016ccg is found to prefer the setup with the JIMWLK evolution included. Our main observation is that final state bulk observables are sensitive to the details of the small-$x$ evolution in the initial state.
@@
To further quantify the differences between the two energy evolution schemes, we calculate the $v_{2}\{4\}/v_{2}\{2\}$ ratio in LHC kinematics. The obtained ratios as a function of centrality are compared to the ALICE data ALICE:2016ccg in Fig. 4. Here, only the “+JIMWLK” setup results in an approximately constant ratio, consistent with the ALICE data. On the other hand, an increasing trend with centrality, especially around small centrality, is obtained in the $Q_{s}(x)$ setup. This qualitative difference can be explained as follows. Because $v_{2}\{2\}$ increases with flow fluctuations, while $v_{2}\{4\}$ decreases (which has been shown explicitly e.g. for Gaussian fluctuations), this ratio is a measure of fluctuations, with smaller values implying larger fluctuations Voloshin:2008dg. As the JIMWLK evolution leads to smoother nuclei, initial state fluctuations are suppressed, leading to a larger ratio than in the $Q_{s}(x)$ setup.
@@
Results for the average transverse momentum at different energies were already shown in Mantysaari:2025tcg. Similarly to the anisotropic flow coefficients, the smoothing (and growth) of the geometry in the +JIMWLK setup led to reduced $\langle p_{T}\rangle$ in the Pb+Pb system at 5.02 TeV and an improved description of the ALICE data ALICE:2018hza.
@@ IV.2 O+O and Ne+Ne collisions
Figure 5: Model predictions for the charged hadron multiplicity distributions in O+O and Ne+Ne collisions at the LHC 5.36 TeV and at RHIC 200 GeV and 71 GeV energies. The preliminary ALICE data is from ALICE:MartaUrioni_IS2025
Oxygen-oxygen collisions provide a unique system to probe the small-$x$ evolution effects in the initial state, because they have been measured at both RHIC at $\sqrt{s}=200$ GeV and LHC at $\sqrt{s}=5.36$ TeV Brewer:2021kiv; ALICE:2025luc; ATLAS:2025nnt; cms:OOnote; Yan:2025kwz. Furthermore, because of the smaller size of the oxygen nuclei compared to lead, initial state fluctuations (and their possible modification due to JIMWLK evolution) can be expected to have a more pronounced effect on flow observables. The LHC also measured Ne+Ne collisions at the same energy as O+O, which is particularly interesting because of the expected differences in shape of these two nuclei Giacalone:2025vxa; cms:OOnote; ALICE:2025luc; ATLAS:2025nnt. The energy dependence in the Ne+Ne system could further be explored by comparing fixed target Ne+Ne collisions (using SMOGII lhcb_smog at LHCb) with Ne+Ne collisions in collider mode.
@@
Figure 6: Collective flow $v_{n}$ in O+O and d+Au collisions at $\sqrt{s}=200$ GeV calculated with the “+JIMWLK” (upper panel) and $Q_{s}(x)$” (lower panel) setups. The preliminary STAR data are from Yan:2025kwz; STAR:2025ivi.
Figure 7: Model predictions for the collective flow $v_{n}\{2\}$, $n=2,3$ (upper panel) and $v_{2}\{4\}$ (lower panel) in O+O and Ne+Ne collisions at 5.36 TeV. The preliminary ALICE data are from ALICE:2025luc.
Figure 8: Model predictions for the collective flow $v_{n}\{2\}$ ratio ($n=2,3$) between O+O and Ne+Ne collisions at 5.36 TeV. The preliminary ALICE data are from ALICE:2025luc.
@@
Comparisons to the RHIC elliptic and triangular flow measurements in Oxygen-Oxygen collisions are shown in Fig. 6. As a reference, the corresponding d+Au measurements at the same center-of-mass energy are also shown. For both initial state prescriptions, two sets of results are shown as upper and lower ends of a band: We either initialize with the full shear-stress tensor $\Pi^{\mu\nu}$ of the energy-momentum tensor extracted from the CYM evolution (our standard method), or set the bulk stress to zero. The bulk stress term does not reflect a real bulk viscosity, but it absorbs differences between the conformal equation of state (EOS) in the Glasma and the QCD EOS in the hydrodynamic medium Schenke:2020mbo. Thus, the band reflects the systematic uncertainty from matching the conformal glasma initial state to the near-equilibrium QCD fluid. We did not show this uncertainty band for observables in heavy ion (e.g. Au, Pb) collisions, because in such large systems the uncertainty from varying the initial viscous tensor is much smaller.
@@
Finally, we study an even smaller collision system, proton-lead collisions at $\sqrt{s}=5.02$ TeV measured at the LHC. In such small systems, the sensitivity to the nucleon geometry is maximal, as the proton event-by-event structure dominates the initial state eccentricities. As already discussed in Ref. Mantysaari:2017cni, proton substructure fluctuations constrained by HERA data are crucial to understand flow harmonics measured in p+Pb collisions (see also an earlier analysis of Ref. Schenke:2014zha). In the companion letter Mantysaari:2025tcg we demonstrated that the multiplicity distributions in such small systems are sensitive to the high-energy evolution in the initial state, and that the ALICE and PHENIX data PHENIX:2003iij; ALICE:2013wgn are better described by the “+JIMWLK” setup. In this work, we focus on flow observables.
@@
Figure 10: Collective flow $v_{n}$ with “+JIMWLK” and $Q_{s}(x)$” setups in p+Pb collisions at the LHC 5.02 TeV. Experimental data is from the ALICE Collaboration ALICE:2019zfl.
@@
The flow harmonics in p+Pb collisions are shown in Fig. 10. We find a larger effect of the inclusion of JIMWLK evolution on $v_{n}\{2\}$ coefficients compared to Pb+Pb collisions shown in Fig. 3. The JIMWLK evolution again decreases the $v_{2}\{2\}$ and $v_{3}\{2\}$ flow harmonics. Unlike in other collision systems studied in this work, the available ALICE data ALICE:2019zfl for proton-nucleus collisions favor the $Q_{s}(x)$ setup. This indicates a preference for stronger subnucleonic fluctuations than those present in the +JIMWLK setup after the evolution. We note, however, that the proton-nucleus collisions are the most challenging systems to describe within a hydrodynamical framework Gallmeister:2018mcn; Schenke:2020mbo; Noronha:2024dtq, and a fully three-dimensional setup may be required to accurately describe correlation measurements involving a rapidity gap in a system where boost invariance is not as good an approximation as in heavy ion collisions.
Figure 11: The $v_{2}^{2}-p_{T}$ correlator in p+Pb collisions at the LHC 5.02 TeV. The ATLAS data is from ATLAS:2019pvn
@@
Finally, we calculate the correlation between the squared elliptic flow $v_{2}^{2}$ and the mean transverse momentum in p+Pb collisions (results for the mean $p_{T}$ were shown in Mantysaari:2025tcg), and present the results in Fig. 11. This correlation is found to depend strongly on the initial-state description. In particular, including the JIMWLK evolution decreases the magnitude of the correlation by $\sim 40\%$ in the central $N_{\mathrm{ch}}\gtrsim 100$ collisions. This sensitivity arises because both the elliptic flow and $\langle p_{T}\rangle$ are directly sensitive to the lumpiness of the nucleon density profile in p+Pb collisions. While including the JIMWLK evolution results in a better description of the ATLAS data ATLAS:2019pvn, the strength of this correlator is still significantly overestimated except in the lowest $N_{\mathrm{ch}}$ bin.

### v2

@@ III Space-time evolution
As mentioned above, the gluon fields after the collision are obtained in the standard way Krasnitz and Venugopalan (1999, 2000, 2001); Schenke et al. (2012) from the Wilson lines determined with one of the above prescriptions. Further, the early-time evolution in the Glasma phase up to $\tau_{0}=0.4$ fm/c is obtained by solving the source-free Yang-Mills equations as implemented in the IP-Glasma model. The energy-momentum tensor $T^{\mu\nu}$ of the system is computed from the gluon fields at $\tau=\tau_{0}$, providing the initial condition for relativistic viscous hydrodynamic simulations, which are performed using MUSIC Schenke et al. (2010, 2011); Paquet et al. (2016).
@@
Since the strong coupling constant factors out of the classical Yang-Mills evolution equations (see discussion in Schenke et al. (2020)), we determine an overall normalization factor $K$ for the energy-momentum tensor $T^{\mu\nu}$ by matching to the measured charged hadron multiplicity in the 0–5% most central Au+Au collisions at $\sqrt{s_{\mathrm{NN}}}=200$ GeV. Once fixed, the JIMWLK evolution predicts the energy dependence of particle production across different collision energies. This stands in contrast to the conventional Glauber-based approach, where the normalization is typically re-tuned separately for each center-of-mass energy. We note that running coupling effects can be expected to result in $K$ having a logarithmic dependence on the center-of-mass energy. As running coupling corrections are included in the JIMWLK evolution employed in this work, analogous corrections could, in principle, be applied to the $K$ factor. This would introduce an additional dependence on the choice of running coupling scheme and is therefore left for future work. Note that $K$ in Table. 1 should not be of order 1 but of order $1/g^{2}$, with the coupling $g$ defined in Eq. (1). Since the classical calculation of the early-time dynamics is independent of $g$, we set $g=1$ in the numerical simulations. Then, $K$ reintroduces the correct dependence of the initial energy density on $\alpha_{s}$. The preferred value $K\simeq 0.1$ corresponds to an effective coupling $g\approx 3$. See the discussion in Ref. Schenke et al. (2020) for more details on this subject. We further remark that in principle $K$ is expected to depend on the collision energy as it is scale dependent.
@@
Focusing exclusively on midrapidity observables, we assume boost invariance in the hydrodynamic evolution. We employ an equation of state derived from lattice QCD calculations Bazavov and others (2014); Moreland and Soltz (2016), and include both shear and bulk viscous effects. These are treated using the Denicol-Niemi-Molnar-Rischke (DNMR) second-order viscous hydrodynamic formalism Denicol et al. (2012), which incorporates spatial gradient corrections up to second order. We do not incorporate an intermediate pre-equilibrium stage. This could be done most directly using the existing KoMPoST framework Kurkela et al. (2019), however there are reasons this is not easily achievable: 1) KoMPoST suffers from numerical instabilities in the presence of large pressure gradients, especially in peripheral and small collision systems. 2) The dynamics of KoMPoST remain conformal, which leads to a longer conformal pre-equilibrium evolution compared to the Glasma phase alone. This results in stronger effects on final-state observables from the discontinuity introduced by switching from the conformal system to one described by the lattice QCD equation of state.
@@
The transition from the fluid to hadronic degrees of freedom is performed on an isothermal hypersurface defined by a constant energy density $e_{\mathrm{sw}}=0.18~\mathrm{GeV/fm}^{3}$, using the Cooper-Frye particlization procedure with viscous corrections implemented via Grad’s 14-moment method Huovinen and Petersen (2012); Shen et al. (2016); Zhao et al. (2023). The resulting hadrons are subsequently evolved in the dilute hadronic phase using the UrQMD transport model Bass and others (1998); Bleicher and others (1999), which accounts for hadronic scatterings and decays.
@@
Here $s$ and $T$ are the entropy density and temperature, respectively. We tune the parameters to reproduce experimental data on charged hadron multiplicity Adler and others (2004), mean transverse momentum Abelev and others (2009), and azimuthal flow coefficients Adamczyk and others (2015) measured in Au+Au collisions at $\sqrt{s_{\mathrm{NN}}}=200$ GeV. The IP-Glasma+JIMWLK setup is used for this tune. The purpose of this work is to quantify the effects of including the small-$x$ JIMWLK evolution in the initial state description. Consequently, the transport properties are constrained using the RHIC data only, as there is only moderate evolution from initial $x_{0}=0.01$ to RHIC kinematics corresponding to $x=0.0025$. The obtained parameters are $\eta_{0}=0.14$, $T_{0}=0.18$ GeV, b=4b=-4 GeV-1, $(\zeta/s)_{\max}=0.123$, $\lambda_{\zeta}=-0.12$, and $w_{\zeta}=0.05$ GeV.
@@ IV.1 Heavy ion collisions
Figure 1: Charged hadron multiplicity distributions in Pb+Pb collisions at the LHC 5.02 TeV and 5.36 TeV, and Au+Au collisions at RHIC 200 GeV energies. The experimental data are taken from Adam and others (2016b); Acharya and others (2025); Adler and others (2005).
@@
Both prescriptions for the initial state energy dependence reproduce the multiplicity in the most central collisions. On the other hand, in more peripheral Pb+Pb collisions beyond a centrality of $\sim 20\%$, inclusion of the JIMWLK evolution results in significantly flatter multiplicity distributions. We interpret this to be due to the geometry evolution included in the “+JIMWLK” setup, which modifies the overlap by blurring the edges of the incoming nuclei, and consequently changes the shape of the multiplicity distribution with centrality. This geometry evolution has been previously studied e.g. in Refs. Schlichting and Schenke (2014); Mäntysaari and Schenke (2018); Mäntysaari et al. (2022a). We have checked that the differences in the viscous entropy production during the late-time hydrodynamic phase are not significant.
@@
Figure 2: Anisotropic flow $v_{n}$ in Au+Au collisions 200 GeV. The experimental data are from Adamczyk and others (2016, 2018); Adam and others (2019).
@@
Figure 3: Collective flow $v_{n}$ in Pb + Pb collisions 5.02 TeV. The experimental data are from Adam and others (2016a).
@@
Figure 4: The $v_{2}\{4\}/v_{2}\{2\}$ ratio in Pb + Pb collisions at 5.02 TeV. The experimental data are from Adam and others (2016a).
@@
The effect of JIMWLK evolution on anisotropic flow coefficients was not studied in the companion letter Mäntysaari et al. (2025b). To set the stage, we first show our results for $v_{2}\{2\},v_{3}\{2\}$ and $v_{4}\{2\}$ in Au+Au collisions at 200 GeV center-of-mass energy in Fig. 2. Note that shear and bulk viscosities have been adjusted to describe this data in the “+JIMWLK” setup, and kept the same in the $Q_{s}(x)$ setup - see discussion in Sec. III. There is only moderate evolution from the initial $x=0.01$ to the typical midrapidity value at RHIC, $x=0.0025$. Yet, there is some difference between our results for $v_{2}\{2\}$ and $v_{2}\{4\}$ for centralities larger than 40% as well as for $v_{3}\{2\}$ for all centralities. In all cases the +JIMWLK result is lower than the corresponding $Q_{s}(x)$ result. This can again be explained by the smoothing effect of the JIMWLK evolution on the initial state geometry, which reduces the initial eccentricities, as we have checked explicitly.
@@
Similar trends are seen for Pb+Pb collisions at 5.02 TeV: The centrality dependence of the flow harmonics $v_{2}\{2\},v_{3}\{2\}$ and $v_{4}\{2\}$ is shown in Fig. 3. Again, the JIMWLK evolution is found to suppress all flow harmonics compared to the $Q_{s}(x)$ setup. Interestingly, the differences are similar to those observed in the lower energy Au+Au collisions, suggesting that the dominant effect from the small-$x$ evolution is the smoothening of the initially lumpy nuclear substructure early in the evolution. Faster evolution at larger $x$ is expected as saturation effects that slow down the evolution become more prominent as $x$ decreases. The ALICE data Adam and others (2016a) is found to prefer the setup with the JIMWLK evolution included. Our main observation is that final state bulk observables are sensitive to the details of the small-$x$ evolution in the initial state.
@@
To further quantify the differences between the two energy evolution schemes, we calculate the $v_{2}\{4\}/v_{2}\{2\}$ ratio in LHC kinematics. The obtained ratios as a function of centrality are compared to the ALICE data Adam and others (2016a) in Fig. 4. Here, only the “+JIMWLK” setup results in an approximately constant ratio, consistent with the ALICE data. On the other hand, an increasing trend with centrality, especially around small centrality, is obtained in the $Q_{s}(x)$ setup. This qualitative difference can be explained as follows. Because $v_{2}\{2\}$ increases with flow fluctuations, while $v_{2}\{4\}$ decreases (which has been shown explicitly e.g. for Gaussian fluctuations), this ratio is a measure of fluctuations, with smaller values implying larger fluctuations Voloshin et al. (2010). As the JIMWLK evolution leads to smoother nuclei, initial state fluctuations are suppressed, leading to a larger ratio than in the $Q_{s}(x)$ setup.
@@
Results for the average transverse momentum at different energies were already shown in Mäntysaari et al. (2025b). Similarly to the anisotropic flow coefficients, the smoothing (and growth) of the geometry in the +JIMWLK setup led to reduced $\langle p_{T}\rangle$ in the Pb+Pb system at 5.02 TeV and an improved description of the ALICE data Acharya and others (2019c).
@@ IV.2 O+O and Ne+Ne collisions
Figure 5: Model predictions for the charged hadron multiplicity distributions in O+O and Ne+Ne collisions at the LHC 5.36 TeV and at RHIC 200 GeV and 71 GeV energies. The preliminary ALICE data is from Urioni
Oxygen-oxygen collisions provide a unique system to probe the small-$x$ evolution effects in the initial state, because they have been measured at both RHIC at $\sqrt{s}=200$ GeV and LHC at $\sqrt{s}=5.36$ TeV J. Brewer, A. Mazeliauskas, and W. van der Schee (2021); I. J. Abualrob et al. (2025); G. Aad et al. (2025); 77; Z. Yan (2025). Furthermore, because of the smaller size of the oxygen nuclei compared to lead, initial state fluctuations (and their possible modification due to JIMWLK evolution) can be expected to have a more pronounced effect on flow observables. The LHC also measured Ne+Ne collisions at the same energy as O+O, which is particularly interesting because of the expected differences in shape of these two nuclei G. Giacalone et al. (2025c); 77; I. J. Abualrob et al. (2025); G. Aad et al. (2025). The energy dependence in the Ne+Ne system could further be explored by comparing fixed target Ne+Ne collisions (using SMOGII 64 at LHCb) with Ne+Ne collisions in collider mode.
@@
Figure 6: Collective flow $v_{n}$ in O+O and d+Au collisions at $\sqrt{s}=200$ GeV calculated with the “+JIMWLK” (upper panel) and $Q_{s}(x)$” (lower panel) setups. The preliminary STAR data are from Z. Yan (2025); 35.
Figure 7: Model predictions for the collective flow $v_{n}\{2\}$, $n=2,3$ (upper panel) and $v_{2}\{4\}$ (lower panel) in O+O and Ne+Ne collisions at 5.36 TeV. The preliminary ALICE data are from Abualrob and others (2025).
Figure 8: Model predictions for the collective flow $v_{n}\{2\}$ ratio ($n=2,3$) between O+O and Ne+Ne collisions at 5.36 TeV. The preliminary ALICE data are from Abualrob and others (2025).
@@
Comparisons to the RHIC elliptic and triangular flow measurements in Oxygen-Oxygen collisions are shown in Fig. 6. As a reference, the corresponding d+Au measurements at the same center-of-mass energy are also shown. For both initial state prescriptions, two sets of results are shown as upper and lower ends of a band: We either initialize with the full shear-stress tensor $\Pi^{\mu\nu}$ of the energy-momentum tensor extracted from the CYM evolution (our standard method), or set the bulk stress to zero. The bulk stress term does not reflect a real bulk viscosity, but it absorbs differences between the conformal equation of state (EOS) in the Glasma and the QCD EOS in the hydrodynamic medium Schenke et al. (2020). Thus, the band reflects the systematic uncertainty from matching the conformal glasma initial state to the near-equilibrium QCD fluid. We did not show this uncertainty band for observables in heavy ion (e.g. Au, Pb) collisions, because in such large systems the uncertainty from varying the initial viscous tensor is much smaller.
@@
Finally, we study an even smaller collision system, proton-lead collisions at $\sqrt{s}=5.02$ TeV measured at the LHC. In such small systems, the sensitivity to the nucleon geometry is maximal, as the proton event-by-event structure dominates the initial state eccentricities. As already discussed in Ref. Mäntysaari et al. (2017), proton substructure fluctuations constrained by HERA data are crucial to understand flow harmonics measured in p+Pb collisions (see also an earlier analysis of Ref. Schenke and Venugopalan (2014)). In the companion letter Mäntysaari et al. (2025b) we demonstrated that the multiplicity distributions in such small systems are sensitive to the high-energy evolution in the initial state, and that the ALICE and PHENIX data Adler and others (2004); Abelev and others (2014b) are better described by the “+JIMWLK” setup. In this work, we focus on flow observables.
@@
Figure 10: Collective flow $v_{n}$ with “+JIMWLK” and $Q_{s}(x)$” setups in p+Pb collisions at the LHC 5.02 TeV. Experimental data is from the ALICE Collaboration Acharya and others (2019b).
@@
The flow harmonics in p+Pb collisions are shown in Fig. 10. We find a larger effect of the inclusion of JIMWLK evolution on $v_{n}\{2\}$ coefficients compared to Pb+Pb collisions shown in Fig. 3. The JIMWLK evolution again decreases the $v_{2}\{2\}$ and $v_{3}\{2\}$ flow harmonics. Unlike in other collision systems studied in this work, the available ALICE data Acharya and others (2019b) for proton-nucleus collisions favor the $Q_{s}(x)$ setup. This indicates a preference for stronger subnucleonic fluctuations than those present in the +JIMWLK setup after the evolution. We note, however, that the proton-nucleus collisions are the most challenging systems to describe within a hydrodynamical framework Gallmeister et al. (2018); Schenke et al. (2020); Noronha et al. (2024), and a fully three-dimensional setup may be required to accurately describe correlation measurements involving a rapidity gap in a system where boost invariance is not as good an approximation as in heavy ion collisions.
Figure 11: The $v_{2}^{2}-p_{T}$ correlator in p+Pb collisions at the LHC 5.02 TeV. The ATLAS data is from Aad and others (2019)
@@
Finally, we calculate the correlation between the squared elliptic flow $v_{2}^{2}$ and the mean transverse momentum in p+Pb collisions (results for the mean $p_{T}$ were shown in Mäntysaari et al. (2025b)), and present the results in Fig. 11. This correlation is found to depend strongly on the initial-state description. In particular, including the JIMWLK evolution decreases the magnitude of the correlation by $\sim 40\%$ in the central $N_{\mathrm{ch}}\gtrsim 100$ collisions. This sensitivity arises because both the elliptic flow and $\langle p_{T}\rangle$ are directly sensitive to the lumpiness of the nucleon density profile in p+Pb collisions. While including the JIMWLK evolution results in a better description of the ATLAS data Aad and others (2019), the strength of this correlator is still significantly overestimated except in the lowest $N_{\mathrm{ch}}$ bin.
In this paper, We simulated 5K minimum-bias events for Pb–Pb and Au–Au collisions, 40K minimum-bias events for p–Pb collision and Ne-Ne at 72 GeV, and 160K for Ne-Ne and O-O at 5.36 TeV. In several figures (Figures. 6-10), the error bands shown represent the systematic uncertainties from matching the Glasma stage to hydrodynamics. The statistical errors are much smaller than these systematic errors and are not explicitly shown.